OpenStudy (anonymous):

The function F(x)=log5 x is increasing. True or false?

jimthompson5910 (jim_thompson5910):

use a graphing calculator. If you don't have one, then use something like desmos https://www.desmos.com/calculator/l7gildqroh

OpenStudy (anonymous):

But what would be the equation? i don't understand it.

jimthompson5910 (jim_thompson5910):

I used the change of base formula \[\Large \log_b(x) = \frac{\log(x)}{\log(b)}\] \[\Large \log_5(x) = \frac{\log(x)}{\log(5)}\]

jimthompson5910 (jim_thompson5910):

the equation you type in is `log(x)/log(5)`

OpenStudy (anonymous):

So it would be true then.

jimthompson5910 (jim_thompson5910):

yep, as you move from left to right, you are moving up on the y axis

OpenStudy (anonymous):

Thank you.

jimthompson5910 (jim_thompson5910):

np

OpenStudy (yumyum247):

hold up i have i quick question here...so what's wrong if we just plot f(x) = log_5 x

OpenStudy (yumyum247):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I don't understand what you're asking @YumYum247

OpenStudy (yumyum247):

why are they both different??

jimthompson5910 (jim_thompson5910):

the plot is shown at the link I posted

jimthompson5910 (jim_thompson5910):

desmos understands `log_5(x)` to mean the same thing

jimthompson5910 (jim_thompson5910):

some calculators won't be able to understand that notation though

OpenStudy (yumyum247):

yes but then i copy pasted the expression above right into the graph and i got a slightly different graph...

jimthompson5910 (jim_thompson5910):

it should be the same

jimthompson5910 (jim_thompson5910):

https://www.desmos.com/calculator/wyw30w55pc

OpenStudy (yumyum247):

hold up let me upload real quick :")

OpenStudy (yumyum247):

http://postimg.org/image/qihrnqonh/

jimthompson5910 (jim_thompson5910):

you typed in log(5x) instead of log_5(x)

OpenStudy (yumyum247):

oh snappp...thanks blurr luv you <3

jimthompson5910 (jim_thompson5910):

https://www.desmos.com/calculator/sp1lzubfxs

OpenStudy (yumyum247):

i got it what chu mean...thanks blurr <3 :")

jimthompson5910 (jim_thompson5910):

no problem